UNIVERSITY OF CALIFORNIA, BERKELEY

BERKELEY · DAVIS · IRVINE · LOS ANGELES · RIVERSIDE · SAN DIEGO · SAN FRANCISCO

�

SANTA BARBARA · SANTA CRUZ

Please Reply To: BERKELEY, CALIFORNIA 94720

Richard J. Fateman Computer Science Division 789 Soda Hall University of California Berkeley, CA 94720-1776 USA

(510) 642-1879

fateman@CS.Berkeley.EDU

July 17, 1997

Fellowship Office Stanford University

To whom it may concern:

This is a letter in reference to Lawrence K. Lo who has applied for a graduate fellowship.

I have known Larry since January 1997 when he served as a teaching assistant in the Spring, 1997 semester for CS61a Structure and Interpretation of Computer Programs, at UC Berkeley. This is a large undergraduate course (350 students) in which I lectured, but the teaching assistants conducted laboratory sessions and discussion sections. Although Larry was an undergraduate at the time, he performed as well as or better than the graduate students assigned to the same task. He did more than his share in preparing exam material, conducting review sessions, and making suggestions for the lectures. He was popular with the students, as well.

I believe he is deserving of financial aid in his graduate studies in computer science.

Sincerely,

Richard J. Fateman

Professor of Computer Science